# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ C.3 Proof of Theorem 3
as $n_{i_{0}}\rightarrow\infty$, under the assumption that $\frac{n_{i_{0}}q_{i_{0}}\mu_{i_{0}}}{\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}}=\omega(\log(q_{i_{0}}n_{i_{0}}))$ provided by Eq. 3.10. Using the same argument we also have
$\lim_{n_{i_{L}}\rightarrow\infty}\sum_{k_{2}=1}^{n_{i_{L}}}\mathbb{E}\big{[}|\tilde{\mathbf{Y}}^{(i_{L})}_{k_{2}}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}^{(i_{L})}_{k_{2}}|>\epsilon\}\big{]}=0.$ (C.21)
By Eq. (C.20), Eq. (C.21), and applying the Lindeberg-Feller central limit theorem (see e.g., Proposition 2.27 in Van der Vaart (2000)) we have
$\tilde{\sigma}_{s,t}^{-1}\Big{[}\sum_{k_{1}=1}^{n_{i_{0}}}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}+\sum_{k_{2}=1}^{n_{i_{L}}}\mathbf{E}^{({i_{L}})}_{t,k_{2}}\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s}\Big{]}\rightsquigarrow\mathcal{N}(0,1)$ (C.22)
as $\min\{n_{i_{0}},n_{i_{L}}\}\rightarrow\infty$. Finally, invoking Theorem 2 and the assumption in Eq. (3.11), we have $\tilde{\sigma}_{s,t}^{-1}(\|\mathbf{R}^{(i_{0},\dots,i_{L})}\|_{\max}+\|\mathbf{S}^{(i_{0},\dots,i_{L})}\|_{\max})\rightarrow 0$ in probability. Then applying Slutsky’s theorem we obtain $\tilde{\sigma}_{s,t}^{-1}(\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i,L}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U_{i_{L}}}})_{s,t}\rightsquigarrow\mathcal{N}(0,1)$ as claimed.
Appendix D Technical Lemmas
D.1 Technical lemmas for Lemma C.1
Lemma D.1.
Consider the setting of Lemma C.1. Then for any $i$, for $\mathbf{E}^{(i)}=\mathbf{A}^{(i)}-\mathbf{P}^{(i)}$ we have
@@ Appendix D Technical Lemmas
$\mathbf{E}^{(i)}=(\mathbf{P}^{(i)}+\mathbf{N}^{(i)})\circ\mathbf{\Omega}^{(i)}/q_{i}-\mathbf{P}^{(i)}=\underbrace{(\mathbf{P}^{(i)}\circ\mathbf{\Omega}^{(i)}/q_{i}-\mathbf{P}^{(i)})}_{\mathbf{E}^{(i,1)}}+\underbrace{\mathbf{N}^{(i)}\circ\mathbf{\Omega}^{(i)}/q_{i}}_{\mathbf{E}^{(i,2)}}.$ (D.1)
@@
$\|\mathbf{E}^{(i,1)}\|$ $\lesssim(n_{i}/q_{i})^{1/2}\|\mathbf{P}^{(i)}\|_{\max}$ (D.2)
@@
$\mathbf{Z}^{(i;s,t)}:=\left\{\begin{aligned} &\mathbf{E}^{(i,1)}_{s,t}\big{[}\mathbf{u}^{(i)}_{s}(\mathbf{u}^{(i)}_{t})^{\top}+\mathbf{u}^{(i)}_{t}(\mathbf{u}^{(i)}_{s})^{\top}\big{]}\text{ for }s<t,\\ &\mathbf{E}^{(i,1)}_{s,s}\mathbf{u}^{(i)}_{s}(\mathbf{u}^{(i)}_{s})^{\top}\qquad\qquad\qquad\quad\text{for }s=t,\end{aligned}\right.$
@@
[…]
@@
$\|\mathbf{U}^{(i)\top}\mathbf{E}^{(i,1)}\mathbf{U}^{(i)}\|$ $\lesssim d^{1/2}q_{i}^{-1/2}\|\mathbf{P}^{(i)}\|_{\max}\log^{1/2}{n_{i}}+dn_{i}^{-1}q_{i}^{-1}\|\mathbf{P}^{(i)}\|_{\max}\log n_{i}$ (D.3)
@@
[…]
@@
$\max\Big{\{}\Big{\|}\sum_{t\in[n_{i}]}\mathbb{E}[(\mathbf{E}^{(i,1)}_{s,t})^{2}\mathbf{u}^{(i)}_{t}(\mathbf{u}^{(i)}_{t})^{\top}]\Big{\|},\Big{\|}\sum_{t\in[n_{i}]}\mathbb{E}[(\mathbf{E}^{(i,1)}_{s,t})^{2}(\mathbf{u}^{(i)}_{t})^{\top}\mathbf{u}^{(i)}_{t}]\Big{\|}\Big{\}}$ $\leq\sigma_{*}^{2}\max\Bigl{\{}\bigl{\|}\mathbf{U}^{(i)}\bigr{\|}^{2},\sum_{t\in[n_{i}]}\|\bm{u}^{(i)}_{t}\|^{2}\Bigr{\}}$
@@
$\|\mathbf{E}^{(i,1)}\mathbf{U}^{(i)}\|_{2\to\infty}\lesssim d^{1/2}q_{i}^{-1/2}\|\mathbf{P}^{(i)}\|_{\max}\log^{1/2}n_{i}$ (D.4)
@@
$\|\mathbf{E}^{(i,2)}\|\lesssim(n_{i}/q_{i})^{1/2}\sigma_{i}$ (D.5)
@@
$\|\mathbf{U}^{(i)\top}\mathbf{E}^{(i,2)}\mathbf{U}^{(i)}\|$ $\lesssim d^{1/2}q_{i}^{-1/2}\sigma_{i}\log^{1/2}n_{i}+dn_{i}^{-1}q_{i}^{-1}\sigma_{i}\log^{3/2}n_{i}$ (D.6)
@@
$\|\mathbf{E}^{(i,2)}\mathbf{U}^{(i)}\|_{2\to\infty}$ $\lesssim d^{1/2}q_{i}^{-1/2}\sigma_{i}\log^{1/2}n_{i}+d^{1/2}n_{i}^{-1/2}q_{i}^{-1}\sigma_{i}\log^{3/2}n_{i}$ (D.7)
@@
Finally we combine Eq. (D.2) and Eq. (D.5), Eq. (D.3) and Eq. (D.6), Eq. (D.4) and Eq. (D.7) and obtain the desired results for $\mathbf{E}^{(i)}$, $\mathbf{U}^{(i)\top}\mathbf{E}^{(i)}\mathbf{U}^{(i)}$, and $\mathbf{E}^{(i)}\mathbf{U}$. ∎
Lemma D.2.
Consider the setting of Lemma C.1. Then for any $i\in[K]$ we have
@@
with high probability, and for $\mathbf{U}^{(i)}$ and $\hat{\mathbf{U}}^{(i)}$ we have
$\|\sin\Theta(\hat{\mathbf{U}}^{(i)},\mathbf{U}^{(i)})\|\leq\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}},$
$\|\mathbf{U}^{(i)\top}\hat{\mathbf{U}}^{(i)}-\mathbf{W}^{(i)\top}_{\mathbf{U}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}},$
$\|\hat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
By perturbation theorem for singular values (see Problem III.6.13 in Horn and Johnson (2012)) and Lemma D.1, for any $k\in[n_{i}]$ we have
@@
with high probability. Then the condition in Eq. (C.1) yields the stated claim for $\lambda_{k}(\mathbf{A}^{(i)})$. Next, by Wedin’s $\sin\Theta$ Theorem (see e.g., Theorem 4.4 in Chapter 4 of Stewart and Sun (1990)) and Lemma D.1, we have
$\|\sin\Theta(\hat{\mathbf{U}}^{(i)},\mathbf{U}^{(i)})\|\leq\frac{\|\mathbf{E}^{(i)}\|}{\lambda_{d}(\mathbf{A}^{(i)})-\lambda_{d+1}(\mathbf{P}^{(i)})}\lesssim\frac{\|\mathbf{E}^{(i)}\|}{\lambda_{i,\min}}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
$\|\mathbf{U}^{(i)\top}\hat{\mathbf{U}}^{(i)}-\mathbf{W}^{(i)\top}_{\mathbf{U}}\|\leq\|\sin\Theta(\hat{\mathbf{U}}^{(i)},\mathbf{U}^{(i)})\|^{2}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}},$
$\|\hat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)}\|\leq\|\sin\Theta(\hat{\mathbf{U}}^{(i)},\mathbf{U}^{(i)})\|+\|\mathbf{U}^{(i)\top}\hat{\mathbf{U}}^{(i)}-\mathbf{W}^{(i)\top}_{\mathbf{U}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
Lemma D.3.
Consider the setting of Lemma C.1. Then for each $i$, we have
$\|\bm{\Lambda}^{(i)}\mathbf{U}^{(i)\top}\hat{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)\top}\hat{\mathbf{U}}^{(i)}\hat{\bm{\Lambda}}^{(i)}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}}+d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i},$
$\|\bm{\Lambda}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)\top}-\mathbf{W}_{\mathbf{U}}^{(i)\top}\hat{\bm{\Lambda}}^{(i)}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}}+d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i},$
$\|(\hat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}\|\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{3/2}}+\frac{d(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}^{1/2}},$
@@
For $\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}}$, because
$\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}}$ $=\mathbf{U}^{\top}\mathbf{P}\hat{\mathbf{U}}-\mathbf{U}^{\top}\mathbf{A}\hat{\mathbf{U}}=-\mathbf{U}^{\top}\mathbf{E}\hat{\mathbf{U}}=-\mathbf{U}^{\top}\mathbf{E}(\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{U}^{\top}\mathbf{E}\mathbf{U}\mathbf{W}_{\mathbf{U}}^{\top},$
by Lemma D.1 and Lemma D.2 we have
$\|\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}}\|$ $\leq\|\mathbf{E}\|\cdot\|\hat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U}\|+\|\mathbf{U}^{\top}\mathbf{E}\mathbf{U}\|$ (D.8)
@@
For $\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}_{\mathbf{U}}^{\top}\hat{\bm{\Lambda}}$, we notice
$\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}_{\mathbf{U}}^{\top}\hat{\bm{\Lambda}}=\bm{\Lambda}(\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{U}^{\top}\hat{\mathbf{U}})+(\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}})+(\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}^{\top})\hat{\bm{\Lambda}}.$
For the right hand side of the above display, we have bounded the second term in Eq. (D.8). For the first term and the third term, by Lemma D.2 we have
$\|\bm{\Lambda}(\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{U}^{\top}\hat{\mathbf{U}})+(\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}^{\top})\hat{\bm{\Lambda}}\|$ $\leq(\|\bm{\Lambda}\|+\|\hat{\bm{\Lambda}}\|)\cdot\|\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}^{\top}\|$ (D.9)
@@
with high probability. Combining Eq. (D.8) and Eq. (D.9), we have
$\|\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}_{\mathbf{U}}^{\top}\hat{\bm{\Lambda}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}}+d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}$ (D.10)

### v4

@@ D.3 Proof of Theorem 3
as $n_{i_{0}}\rightarrow\infty$, under the assumption that $\frac{n_{i_{0}}q_{i_{0}}\mu_{i_{0}}}{\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}}=\omega(\log(q_{i_{0}}n_{i_{0}}))$ provided by Eq. (3.13). Using the same argument we also have
$\lim_{n_{i_{L}}\rightarrow\infty}\sum_{k_{2}=1}^{n_{i_{L}}}\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}^{(i_{L})}_{k_{2}}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}^{(i_{L})}_{k_{2}}|>\epsilon\}\big{]}=0.$ (D.21)
By Eq. (D.20), Eq. (D.21), and applying the Lindeberg-Feller central limit theorem (see e.g., Proposition 2.27 in Van der Vaart (2000)) we have
$\widetilde{\sigma}_{s,t}^{-1}\Big{[}\sum_{k_{1}=1}^{n_{i_{0}}}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}+\sum_{k_{2}=1}^{n_{i_{L}}}\mathbf{E}^{({i_{L}})}_{t,k_{2}}\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s}\Big{]}\rightsquigarrow\mathcal{N}(0,1)$ (D.22)
as $\min\{n_{i_{0}},n_{i_{L}}\}\rightarrow\infty$. Finally, invoking Theorem 2 and the assumption in Eq. (3.14), we have $\widetilde{\sigma}_{s,t}^{-1}(\|\mathbf{R}^{(i_{0},\dots,i_{L})}\|_{\max}+\|\mathbf{S}^{(i_{0},\dots,i_{L})}\|_{\max})\rightarrow 0$ in probability. Then applying Slutsky’s theorem we obtain $\widetilde{\sigma}_{s,t}^{-1}(\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i,L}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U_{i_{L}}}})_{s,t}\rightsquigarrow\mathcal{N}(0,1)$ as claimed.
Appendix E Technical Lemmas
E.1 Technical lemmas for Lemma D.1
Lemma E.1.
Consider the setting of Lemma D.1. Then for any $i$, for $\mathbf{E}^{(i)}=\mathbf{A}^{(i)}-\mathbf{P}^{(i)}$ we have
@@ Appendix E Technical Lemmas
$\mathbf{E}^{(i)}=(\mathbf{P}^{(i)}+\mathbf{N}^{(i)})\circ\mathbf{\Omega}^{(i)}/q_{i}-\mathbf{P}^{(i)}=\underbrace{(\mathbf{P}^{(i)}\circ\mathbf{\Omega}^{(i)}/q_{i}-\mathbf{P}^{(i)})}_{\mathbf{E}^{(i,1)}}+\underbrace{\mathbf{N}^{(i)}\circ\mathbf{\Omega}^{(i)}/q_{i}}_{\mathbf{E}^{(i,2)}}.$ (E.1)
@@
$\|\mathbf{E}^{(i,1)}\|$ $\lesssim(n_{i}/q_{i})^{1/2}\|\mathbf{P}^{(i)}\|_{\max}$ (E.2)
@@
$\mathbf{Z}^{(i;s,t)}=\left\{\begin{aligned} &\mathbf{E}^{(i,1)}_{s,t}\big{[}\mathbf{u}^{(i)}_{s}(\mathbf{u}^{(i)}_{t})^{\top}+\mathbf{u}^{(i)}_{t}(\mathbf{u}^{(i)}_{s})^{\top}\big{]}\text{ for }s<t,\\ &\mathbf{E}^{(i,1)}_{s,s}\mathbf{u}^{(i)}_{s}(\mathbf{u}^{(i)}_{s})^{\top}\qquad\qquad\qquad\quad\text{for }s=t,\end{aligned}\right$.
@@
Furthermore, as $\mathbb{E}[(\mathbf{E}^{(i,1)}_{s,t})^{2}]=\frac{1-q_{i}}{q_{i}}(\mathbf{P}^{(i)}_{s,t})^{2}\leq q_{i}^{-1}\|\mathbf{P}^{(i)}\|_{\max}^{2}$ for all $s,t\in[n_{i}]$, we have
@@
$\|\mathbf{U}^{(i)\top}\mathbf{E}^{(i,1)}\mathbf{U}^{(i)}\|$ $\lesssim d^{1/2}q_{i}^{-1/2}\|\mathbf{P}^{(i)}\|_{\max}\log^{1/2}{n_{i}}+dn_{i}^{-1}q_{i}^{-1}\|\mathbf{P}^{(i)}\|_{\max}\log n_{i}$ (E.3)
@@
Next note that $\|\mathbf{E}^{(i,1)}\mathbf{U}^{(i)}\|_{2\to\infty}=\max_{s\in[n_{i}]}\|(\mathbf{E}^{(i,1)}\mathbf{U}^{(i)})_{s}\|$, where $(\mathbf{E}^{(i,1)}\mathbf{U}^{(i)})_{s}$ is the $s$th row of $\mathbf{E}^{(i,1)}\mathbf{U}^{(i)}$. Thus, for any fixed $s\in[n_{i}]$, $(\mathbf{E}^{(i,1)}\mathbf{U}^{(i)})_{s}=\sum_{t\in[n_{i}]}[\mathbf{E}^{(i,1)}_{s,t}\mathbf{u}^{(i)}_{t}$]; note that $\{\mathbf{E}^{(i,1)}_{s,t}\mathbf{u}^{(i)}_{t}\}_{t\in[n_{i}]}$ are independent, random matrices of dimension $1\times d$ with $\mathbb{E}[\mathbf{E}^{(i,1)}_{s,t}\mathbf{u}^{(i)}_{t}]=\mathbf{0}$ and
@@
$\max\Big{\{}\Big{\|}\sum_{t\in[n_{i}]}\mathbb{E}[(\mathbf{E}^{(i,1)}_{s,t})^{2}\mathbf{u}^{(i)}_{t}(\mathbf{u}^{(i)}_{t})^{\top}]\Big{\|},\Big{\|}\sum_{t\in[n_{i}]}\mathbb{E}[(\mathbf{E}^{(i,1)}_{s,t})^{2}(\mathbf{u}^{(i)}_{t})^{\top}\mathbf{u}^{(i)}_{t}$$\Big{\|}\Big{\}}$ $\leq\sigma_{*}^{2}\max\Bigl{\{}\bigl{\|}\mathbf{U}^{(i)}\bigr{\|}^{2},\sum_{t\in[n_{i}]}\|\bm{u}^{(i)}_{t}\|^{2}\Bigr{\}}$
@@
$\|\mathbf{E}^{(i,1)}\mathbf{U}^{(i)}\|_{2\to\infty}\lesssim d^{1/2}q_{i}^{-1/2}\|\mathbf{P}^{(i)}\|_{\max}\log^{1/2}n_{i}$ (E.4)
@@
$\|\mathbf{E}^{(i,2)}\|\lesssim(n_{i}/q_{i})^{1/2}\sigma_{i}$ (E.5)
@@
$\|\mathbf{U}^{(i)\top}\mathbf{E}^{(i,2)}\mathbf{U}^{(i)}\|$ $\lesssim d^{1/2}q_{i}^{-1/2}\sigma_{i}\log^{1/2}n_{i}+dn_{i}^{-1}q_{i}^{-1}\sigma_{i}\log^{3/2}n_{i}$ (E.6)
@@
$\|\mathbf{E}^{(i,2)}\mathbf{U}^{(i)}\|_{2\to\infty}$ $\lesssim d^{1/2}q_{i}^{-1/2}\sigma_{i}\log^{1/2}n_{i}+d^{1/2}n_{i}^{-1/2}q_{i}^{-1}\sigma_{i}\log^{3/2}n_{i}$ (E.7)
@@
Finally we combine Eq. (E.2) and Eq. (E.5), Eq. (E.3) and Eq. (E.6), Eq. (E.4) and Eq. (E.7) and obtain the desired results for $\mathbf{E}^{(i)}$, $\mathbf{U}^{(i)\top}\mathbf{E}^{(i)}\mathbf{U}^{(i)}$, and $\mathbf{E}^{(i)}\mathbf{U}$. ∎
Lemma E.2.
Consider the setting of Lemma D.1. Then for any $i\in[K]$ we have
@@
with high probability, and for $\mathbf{U}^{(i)}$ and $\widehat{\mathbf{U}}^{(i)}$ we have
$\|\sin\Theta(\widehat{\mathbf{U}}^{(i)},\mathbf{U}^{(i)})\|\leq\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}},$
$\|\mathbf{U}^{(i)\top}\widehat{\mathbf{U}}^{(i)}-\mathbf{W}^{(i)\top}_{\mathbf{U}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}},$
$\|\widehat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
By perturbation theorem for singular values (see Problem III.6.13 in Horn and Johnson (2012)) and Lemma E.1, for any $k\in[n_{i}]$ we have
@@
with high probability. Then the condition in Eq. (D.1) yields the stated claim for $\lambda_{k}(\mathbf{A}^{(i)})$. Next, by Wedin’s $\sin\Theta$ Theorem (see e.g., Theorem 4.4 in Chapter 4 of Stewart and Sun (1990)) and Lemma E.1, we have
$\|\sin\Theta(\widehat{\mathbf{U}}^{(i)},\mathbf{U}^{(i)})\|\leq\frac{\|\mathbf{E}^{(i)}\|}{\lambda_{d}(\mathbf{A}^{(i)})-\lambda_{d+1}(\mathbf{P}^{(i)})}\lesssim\frac{\|\mathbf{E}^{(i)}\|}{\lambda_{i,\min}}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
$\|\mathbf{U}^{(i)\top}\widehat{\mathbf{U}}^{(i)}-\mathbf{W}^{(i)\top}_{\mathbf{U}}\|\leq\|\sin\Theta(\widehat{\mathbf{U}}^{(i)},\mathbf{U}^{(i)})\|^{2}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{2}},$
$\|\widehat{\mathbf{U}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{U}^{(i)}\|\leq\|\sin\Theta(\widehat{\mathbf{U}}^{(i)},\mathbf{U}^{(i)})\|+\|\mathbf{U}^{(i)\top}\widehat{\mathbf{U}}^{(i)}-\mathbf{W}^{(i)\top}_{\mathbf{U}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}$
@@
Lemma E.3.
Consider the setting of Lemma D.1. Then for each $i$, we have
$\|\bm{\Lambda}^{(i)}\mathbf{U}^{(i)\top}\widehat{\mathbf{U}}^{(i)}-\mathbf{U}^{(i)\top}\widehat{\mathbf{U}}^{(i)}\widehat{\bm{\Lambda}}^{(i)}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}}+d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i},$
$\|\bm{\Lambda}^{(i)}\mathbf{W}_{\mathbf{U}}^{(i)\top}-\mathbf{W}_{\mathbf{U}}^{(i)\top}\widehat{\bm{\Lambda}}^{(i)}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}}+d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i},$
$\|(\widehat{\bm{\Lambda}}^{(i)})^{1/2}\mathbf{W}^{(i)}_{\mathbf{U}}-\mathbf{W}^{(i)}_{\mathbf{U}}(\bm{\Lambda}^{(i)})^{1/2}\|\lesssim\frac{d^{1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}^{3/2}}+\frac{d(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}^{1/2}},$
@@
For $\bm{\Lambda}\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{U}^{\top}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}}$, because
$\bm{\Lambda}\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{U}^{\top}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}}$ $=\mathbf{U}^{\top}\mathbf{P}\widehat{\mathbf{U}}-\mathbf{U}^{\top}\mathbf{A}\widehat{\mathbf{U}}=-\mathbf{U}^{\top}\mathbf{E}\widehat{\mathbf{U}}=-\mathbf{U}^{\top}\mathbf{E}(\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U})\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{U}^{\top}\mathbf{E}\mathbf{U}\mathbf{W}_{\mathbf{U}}^{\top},$
by Lemma E.1 and Lemma E.2 we have
$\|\bm{\Lambda}\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{U}^{\top}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}}\|$ $\leq\|\mathbf{E}\|\cdot\|\widehat{\mathbf{U}}\mathbf{W}_{\mathbf{U}}-\mathbf{U}\|+\|\mathbf{U}^{\top}\mathbf{E}\mathbf{U}\|$ (E.8)
@@
For $\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}_{\mathbf{U}}^{\top}\widehat{\bm{\Lambda}}$, we notice
$\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}_{\mathbf{U}}^{\top}\widehat{\bm{\Lambda}}=\bm{\Lambda}(\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{U}^{\top}\widehat{\mathbf{U}})+(\bm{\Lambda}\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{U}^{\top}\widehat{\mathbf{U}}\widehat{\bm{\Lambda}})+(\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}^{\top})\widehat{\bm{\Lambda}}$.
For the right hand side of the above display, we have bounded the second term in Eq. (E.8). For the first term and the third term, by Lemma E.2 we have
$\|\bm{\Lambda}(\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{U}^{\top}\widehat{\mathbf{U}})+(\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}^{\top})\widehat{\bm{\Lambda}}\|$ $\leq(\|\bm{\Lambda}\|+\|\widehat{\bm{\Lambda}}\|)\cdot\|\mathbf{U}^{\top}\widehat{\mathbf{U}}-\mathbf{W}_{\mathbf{U}}^{\top}\|$ (E.9)
@@
with high probability. Combining Eq. (E.8) and Eq. (E.9), we have
$\|\bm{\Lambda}\mathbf{W}_{\mathbf{U}}^{\top}-\mathbf{W}_{\mathbf{U}}^{\top}\widehat{\bm{\Lambda}}\|\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})^{2}n_{i}}{q_{i}\lambda_{i,\min}}+d^{1/2}q_{i}^{-1/2}(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}$ (E.10)